OpenStudy (ccchristian1):

write an equation in point slope form for the line through the given point with the given slope. (5,2);m=3 a.y+2=3(x-5) b.y+2=3x-5 c.y-2=3(x+5) d.y-2=3(x-5)

OpenStudy (ccchristian1):

my answer was c. i don't know if i am right

OpenStudy (ivettef365):

that is a formula that goes like this y -y1 = m(x-x1) or something like this

OpenStudy (ivettef365):

all you need to do is plug the numbers you have coordinate (5,2) x = 5 y = 2 and m=3

OpenStudy (ivettef365):

is not C

OpenStudy (ivettef365):

do you understand, just plug in the numbers in the formula

OpenStudy (ccchristian1):

a little bit

OpenStudy (ccchristian1):

i plug them in but i get wrong answer

OpenStudy (ivettef365):

y - (2) = 3 (x - (5))

OpenStudy (ccchristian1):

how did get negative ?

OpenStudy (ccchristian1):

not saying u are wrong. i just want to understand

OpenStudy (ivettef365):

the formula is y - y1 = m(x - x1) if you would have a coordinate with a negative number for example (-2,-5) then it would turn into positive y - (-5) = m (x - (-2))

OpenStudy (ivettef365):

then it would have being y +5 = m (x +2) do you see it

OpenStudy (ccchristian1):

so everything just goes opposite ?

OpenStudy (ivettef365):

correct

OpenStudy (ccchristian1):

what happens if its like (-5,2)

OpenStudy (ivettef365):

then it would be y - (2) = m (x - (-5)) which is y - 2 = m (x +5)

OpenStudy (ccchristian1):

wow okay i understand now. thank you!

OpenStudy (ivettef365):

you're welcome :)